# Exploiting Task Relationships for Continual Learning Using Transferability-Aware Task Embeddings

### v2

@@ A.1.1 Rehearsal-Free CL
The strictness of CL settings varies with the extent of allowed previous data accessibility. Multi-task learning [Caruana, 1997], with full data availability of all tasks, can actually be viewed as a special case of CL, while rehearsal-free CL [Smith et al., 2023b], with no previous data involved in the training of new tasks, is the strictest CL setting under this criteria. Despite the success of rehearsal-based methods in various benchmarks [Bang et al., 2021, Shin et al., 2017, Belouadah and Popescu, 2019], rehearsal-free CL is catching the attention of researchers recently [Smith et al., 2023b] because of its low dependency on revisiting previous tasks and therefore broader application in the era of growing data privacy concerns. Existing works on rehearsal-free CL are mostly based on regularization strategies. EWC [Kirkpatrick et al., 2017] and SI [Zenke et al., 2017] introduce penalties to restrict the alteration of parameters vital for addressing prior tasks, thereby reducing the risk of CF. LwF [Li and Hoiem, 2017, Rebuffi et al., 2017] proposes a cross-entropy loss between the predicted class distribution of the (n-1)-th task, as generated by the model before and after learning the n-th task. Smith et al. [2023b] reviews these methods and proposes regularization combinations for better CL performance. In this work, we follow these works and focus on the more challenging rehearsal-free CL setting.
@@ A.1.2 TIL versus CIL
Based on the discrepancy between $D_{j-1}$ and $D_{j}$, Hsu et al. [2018] and Van de Ven and Tolias [2019] categorize CL settings into three specific scenarios: task incremental, class incremental, and domain incremental. Table 3 summarizes the differences among these scenarios. For a better concentration on the study of CL methodology, our work mainly focuses on the task incremental CL (TIL) . In this scenario, the output spaces of tasks are partitioned by task IDs and mutually exclusive between $D_{j-1}$ and $D_{j}$, which is denoted as $\mathbf{Y}^{(j-1)}\neq\mathbf{Y}^{(j)}$. It can be then naturally indicated that $P(\mathbf{Y}^{(j-1)})\neq P(\mathbf{Y}^{(j)})$ and $P(\mathbf{X}^{(j-1)})\neq P(\mathbf{X}^{(j)})$. Notably, here task IDs are accessible during both training and testing. An adaptation of our method to other CL settings could be effectively conducted by introducing additional task inference modules similar to previous works. As an instance, we made a trial in this work that adapts our framework to class incremental CL (CIL) in Sec. 5.3 using a task ID classifier trained on image features.
@@ A.2 HGR maximal correlation for self transferability $H(T_{j},T_{j})$
Here, the correlation is derived by taking the maximum over all functions $f$, $g$ with zero mean and unit variance, which hence extracts the most correlated aspects of $X$ and $Y$. This definition is equivalent to the maximum of the two-sided H-score given by Huang et al. [2019] with normalized functions $f$ and $g$:
@@
Notably, Eqn. 14 is exactly the definition of H-score transferability [Bao et al., 2019], which is actually proposed by applying the H-score theories to transfer learning scenarios. Therefore, with normalized $f$, $g$, the HGR maximal correlation between $X_{j}$, $Y_{j}$ is mathematically equivalent to the H-score metric of the theoretical optimal model $f^{*}_{j}$ on task $j$.
@@
A.3 Algorithm for H-embedding Guided Hypernet Framework
@@
       for $n\leftarrow 1$ to $j-1$ do // Compute transferability
@@
             $H(T_{n},T_{j})\leftarrow tr\left(cov(f_{l}(x^{(j)},\Theta^{(n)}))^{-1}cov(\mathbb{E}_{P_{X|Y}}[f_{l}(x^{(j)},\Theta^{(n)})|y^{(j)}])\right)$triangleright Eq. 2
@@
       $\hat{e}^{(j)},\gamma^{(j)}\leftarrow\operatorname*{arg\,min}_{\hat{e}^{(j)},\gamma^{(j)}}\sum_{n=1}^{j-1}\left(||\hat{e}^{(j)}-e^{(n)}||_{2}-\gamma^{(j)}\exp(-\mathcal{AHP}(T_{n},T_{j}))\right)^{2}$triangleright Eq. 5
@@
       $\Theta_{h}\leftarrow\Theta_{h}-\lambda\nabla_{\Theta_{h}}L$triangleright Eq. 9
@@
A.4 Benchmarks of Experiments
PermutedMNIST [Goodfellow et al., 2013] (N=10) benchmark is a variant of MNIST [LeCun et al., 1998], forming CL tasks from the original MNIST dataset by applying random permutations to the input image pixels. The permuting procedure can be repeated 9 times in experiments to yield a task sequence of 10, with each task consisting of 70,000 images (60,000 for training and 10,000 for testing) of digits from 0 to 9. CIFAR-100 [Krizhevsky et al., 2009] (N = 10) is a benchmark composed of 10 ten-way classification tasks composed by splitting a CIFAR-100 dataset into ten tasks. The model is sequentially trained on the tasks, each with 6,000 images (5,000 for training and 1,000 for testing). ImageNet-R [Hendrycks et al., 2021], built upon the ImageNet dataset [Deng et al., 2009], features a diverse range of renditions of ImageNet classes. This benchmark includes a total of 30,000 images across 200 classes from ImageNet. For continual learning evaluation, ImageNet-R (N = 5, 10, 20) is formed by organizing ImageNet-R into 5, 10, and 20 tasks respectively, each containing 40, 20, and 10 classes and around 6000, 3000, 1500 samples (roughly 5/6 for training and 1/6 for testing). DomainNet [Peng et al., 2019] (N = 5) is a benchmark derived from the 345 classes classification task across six distinct domains, by fusing all domains and splitting into 5 * 69 classification tasks.
A.5 Experimental settings
A.5.1 Comparison Experiments (Table 1)
@@
Our selection of baselines in this work aims to encompass a wide range of baseline categories, covering two of three primary categories in contemporary CL researches (i.e., replay-based, regularization-based, and architecture-based), with the replay-based methods not conforming to our rehearsal-free setting. The specific choice of baselines in each category is mainly based on performance comparison conclusions in recent works such as Smith et al. [2023b] and Kang et al. [2022]. Therefore, we believe that our comparison study has included the most competitive and representative baselines.
@@
In CIFAR-100 and ImageNet-R(10 tasks) datasets with the ResNet-32 backbone network, we evaluate various baseline methods applicable to its model structure including Full Finetune, EWC, L2, PredKD+FeatKD, PackNet, HyperNet, WSN. All baselines are implemented on our own using a ResNet-32 with ‘option A’, i.e., leveraging the zero-padding shortcuts for increasing dimensions. In CIFAR-100 and ImageNet-R datasets with the ViT-B/16 backbone network, we mainly evaluate PEFT baseline methods including L2P, DualPrompt, CODA-Prompt, HiDe-Prompt, InfLoRA, and SD-LoRA. We cited the results of most baselines from previous works [Wang et al., 2023, Wu et al., 2025]. For ResNet, the experiments on CIFAR-100 are conducted on NVIDIA GeForce RTX 3090 GPUs with 100 epochs of training (unless early-stop), and the ImageNet-R experiments are carried out on NVIDIA A800 or A100 GPUs with 200 epochs of training (unless early-stop). For ViT, all experiments are conducted on NVIDIA A800 or A100 GPUs until convergence.
@@
For the EWC baseline [Kirkpatrick et al., 2017], we add a regularization term to the loss function to penalize significant changes to parameters important for previously learned tasks. The importance of each parameter is estimated using the Fisher Information Matrix. The regularization coefficient $\lambda$ is set to 10, following standard practice.
@@
For the PredKD + FeatKD method [Smith et al., 2023b], we incorporate both prediction distillation and feature distillation to transfer knowledge from previous tasks to new ones. The distillation loss combines the Kullback-Leibler divergence between the soft outputs of the teacher (model trained on previous tasks) and the student (current model), as well as the mean squared error between their intermediate feature representations. The loss weights are set to $\alpha=1.0$ and $\beta=0.5$ based on preliminary tuning.
In the PackNet method [Mallya and Lazebnik, 2018], we employ iterative pruning to allocate dedicated network weights for each task. After training on each task, we prune a certain percentage of the weights with the smallest magnitudes. Following the recommendations in the original paper, we experiment with pruning rates of 0.5, 0.75, and 0.8. We select the pruning rate of 0.8, which yields the best performance in our setting. After pruning, we fine-tune the remaining weights for an additional 10 epochs with a reduced learning rate of $1\times 10^{-4}$.
For HyperNet method [von Oswald et al., 2020], we primarily follow the original work in training settings, with the learning rate being 0.001, the CL loss beta being 0.05, and scheduling and transforming strategies being the same as those used by Oswald. The embedding dimension is set to 32.
For the WSN method [Kang et al., 2022], we follow its original paper and use the default values of parameters in its official code repository. We choose the sparsity parameter $c=0.5$ which performs best as listed in the WSN literature. Other parameters are set to the following values: optimization via Adam, a learning rate initialized at 1e-3 with a minimum of 1e-6, and a patience of 6 epochs for reducing the learning rate by a factor of 2. The models are trained for 100 epochs, with a batch size of 64 for both training and testing.
For H-embedding guided hypernet, the learning rate is set to 0.0005, with the embedding loss beta and CL loss beta both set to 0.05. The scheduling and transforming strategies are set the same as those of von Oswald et al. [2020] and the embedding dimension is also set to 32. For the learning of H-embedding, we update Eqn. 5 using gradient descent for 2000 iterations and the $f$, $g$ in HGR maximal correlation for 100 epochs respectively, both using a subset of 1000 samples from the training set.
@@
In all methods, we adhere to the principles of continual learning by not tuning hyperparameters on the full task set. In hypernet generation, special care was taken in handling batch normalization layers, especially in methods involving parameter freezing or pruning. We store and update batch normalization statistics separately for each task to ensure proper normalization during both training and inference. Also, to ensure that the parameter size of the hypernet framework should not exceed that of a normal network, we follow von Oswald et al. [2020] and chunk the weights into size of 7000, using additional chunking embedding to generate each chunked weight.
A.5.2 Additional Comparison Experiments (Table 2)
@@
A.6 Ablation Studies
A.6.1 H-embed Hnet - LoRA (ViT) Ablation
@@
A.6.2 H-embed Hnet - Full Generation Ablation
@@
Considering the smaller data dimension and model size in this setting, the embedding size is reduced to 24 and the training iteration number is set to 5000. The backbone model on PermutedMNIST is selected to be an MLP with fully-connected layers of size 1000, 1000 as used by Van de Ven and Tolias [2019]. We configure the learning rate as 0.0001 and the embedding loss beta as 0.05. The results are derived on NVIDIA GeForce RTX 3090 GPUs.
@@
The tasks in this setting are derived the same as in comparison studies. Yet, the backbone model is differently set to a 4-layer CNN as used by Zenke et al. [2017]. We also follow Oswald in most of the hyperparameters, configuring the learning rate to 0.0001, embedding size to 32, as well as using the same scheduling strategies. We train each method with 100 epochs and the embedding loss beta is set to 0.2 for H-embed and rand-embed hypernets. The results are derived on NVIDIA GeForce RTX 3090 GPUs.
@@
A.7 In-depth Performance Analysis

### v3

@@ A.1.1 Rehearsal-Free CL
The strictness of CL settings varies with the extent of allowed previous data accessibility. Multi-task learning [caruana1997multitask], with full data availability of all tasks, can actually be viewed as a special case of CL, while rehearsal-free CL [smith2023closer], with no previous data involved in the training of new tasks, is the strictest CL setting under this criteria. Despite the success of rehearsal-based methods in various benchmarks [bang2021rainbow, shin2017continual, belouadah2019il2m], rehearsal-free CL is catching the attention of researchers recently [smith2023closer] because of its low dependency on revisiting previous tasks and therefore broader application in the era of growing data privacy concerns. Existing works on rehearsal-free CL are mostly based on regularization strategies. EWC [kirkpatrick2017overcoming] and SI [zenke2017continual] introduce penalties to restrict the alteration of parameters vital for addressing prior tasks, thereby reducing the risk of CF. LwF [li2017learning, rebuffi2017icarl] proposes a cross-entropy loss between the predicted class distribution of the (n-1)-th task, as generated by the model before and after learning the n-th task. smith2023closer reviews these methods and proposes regularization combinations for better CL performance. In this work, we follow these works and focus on the more challenging rehearsal-free CL setting.
@@ A.1.2 TIL versus CIL
Based on the discrepancy between $D_{j-1}$ and $D_{j}$, hsu2018re and van2019three categorize CL settings into three specific scenarios: task incremental, class incremental, and domain incremental. Table 3 summarizes the differences among these scenarios. For a better concentration on the study of CL methodology, our work mainly focuses on the task incremental CL (TIL) . In this scenario, the output spaces of tasks are partitioned by task IDs and mutually exclusive between $D_{j-1}$ and $D_{j}$, which is denoted as $\mathbf{Y}^{(j-1)}\neq\mathbf{Y}^{(j)}$. It can be then naturally indicated that $P(\mathbf{Y}^{(j-1)})\neq P(\mathbf{Y}^{(j)})$ and $P(\mathbf{X}^{(j-1)})\neq P(\mathbf{X}^{(j)})$. Notably, here task IDs are accessible during both training and testing. An adaptation of our method to other CL settings could be effectively conducted by introducing additional task inference modules similar to previous works. As an instance, we made a trial in this work that adapts our framework to class incremental CL (CIL) in Sec. 5.3 using a task ID classifier trained on image features.
@@ A.2 HGR maximal correlation for self transferability $H(T_{j},T_{j})$
Here, the correlation is derived by taking the maximum over all functions $f$, $g$ with zero mean and unit variance, which hence extracts the most correlated aspects of $X$ and $Y$. This definition is equivalent to the maximum of the two-sided H-score given by huang2019information with normalized functions $f$ and $g$:
@@
Notably, Eqn. 14 is exactly the definition of H-score transferability [bao2019information], which is actually proposed by applying the H-score theories to transfer learning scenarios. Therefore, with normalized $f$, $g$, the HGR maximal correlation between $X_{j}$, $Y_{j}$ is mathematically equivalent to the H-score metric of the theoretical optimal model $f^{*}_{j}$ on task $j$.
@@
A.3 Discussion on the Use of H-Score and Alternative Task Embeddings
While technically task embeddings can be estimated using various methods, the alternatives are generally less suitable for our target setting, as they either require access to source data for each task—assumptions that are infeasible under rehearsal-free CL—or do not explicitly model task-to-task relationships, thereby offering limited benefits in facilitating bidirectional transfer across CL tasks.
To address this, we propose a new formulation where task embeddings are computed online by explicitly optimizing them to approximate a transferability metric across tasks. In this design, we adopt H-score as the target transferability metric primarily due to its source-data-free nature, which allows us to maintain compliance with CL constraints. Additionally, H-score is theoretically grounded, computationally efficient, and empirically shown to correlate well with actual transfer performance. These characteristics make it a particularly suitable and representative choice for our embedding algorithm. Yet, our framework is not inherently tied to H-score, and there admittedly exist other source-free embeddings or transferability metrics (e.g. Task2Vec [achille2019task2vec], LEEP [nguyen2020leep], LogME [you2021logme]) that could also, in principle, be used within it.
As both an ablation and a further assessment of our framework’s flexibility, we conducted preliminary experiments on ImageNet-R 10 task benchmark by:
  • Replacing H-score with other source-free transferability metrics (LEEP and LogME) while keeping the embedding algorithm and overall pipeline unchanged.
  • Replacing the entire H-embedding formulation with Task2Vec.
The results are shown in Tab. 4.
Method H-Embedding* H-Embedding (LEEP) H-Embedding (LogME) Task2Vec
FAA (%) 81.8 81.6 81.4 79.8
Table 4: Comparison of different task embedding variants on the ImageNet-R 10-task benchmark.
As shown, replacing the transferability metric alone leads to only minor drop in performance. This supports our view that while H-score is a strong instantiation, the core contribution of our work lies in the transferability-guided embedding formulation and its integration into a hypernetwork-based parameter generation pipeline—a general framework that remains valid even when alternative metrics are employed.
However, replacing the entire embedding formulation with task2vec leads to a more noticeable performance degradation. We believe this is because Task2Vec computes task embeddings independently without explicitly modeling task-to-task relationships, which limits its capacity to guide transfer dynamics in CL. Additionally, Task2Vec is highly sensitive to the choice and training of the probe network, which introduces extra variance in effectiveness and incurs additional memory costs—challenges not present in our H-embedding approach.
A.4 Algorithm for H-embedding Guided Hypernet Framework
@@
 for $n\leftarrow 1$ to j1j-1 do // Compute transferability
@@
    $H(T_{n},T_{j})\leftarrow tr\left(cov(f_{l}(x^{(j)},\Theta^{(n)}))^{-1}cov(\mathbb{E}_{P_{X|Y}}[f_{l}(x^{(j)},\Theta^{(n)})|y^{(j)}])\right)$ $\triangleright$ Eq. 2
@@
[…]
@@
 $\Theta_{h}\leftarrow\Theta_{h}-\lambda\nabla_{\Theta_{h}}L$ $\triangleright$ Eq. 9
@@
A.5 Benchmarks of Experiments
PermutedMNIST [goodfellow2013empirical] (N=10) benchmark is a variant of MNIST [lecun1998gradient], forming CL tasks from the original MNIST dataset by applying random permutations to the input image pixels. The permuting procedure can be repeated 9 times in experiments to yield a task sequence of 10, with each task consisting of 70,000 images (60,000 for training and 10,000 for testing) of digits from 0 to 9. CIFAR-100 [krizhevsky2009learning] (N = 10) is a benchmark composed of 10 ten-way classification tasks composed by splitting a CIFAR-100 dataset into ten tasks. The model is sequentially trained on the tasks, each with 6,000 images (5,000 for training and 1,000 for testing). ImageNet-R [hendrycks2021many], built upon the ImageNet dataset [deng2009imagenet], features a diverse range of renditions of ImageNet classes. This benchmark includes a total of 30,000 images across 200 classes from ImageNet. For continual learning evaluation, ImageNet-R (N = 5, 10, 20) is formed by organizing ImageNet-R into 5, 10, and 20 tasks respectively, each containing 40, 20, and 10 classes and around 6000, 3000, 1500 samples (roughly 5/6 for training and 1/6 for testing). DomainNet [peng2019moment] (N = 5) is a benchmark derived from the 345 classes classification task across six distinct domains, by fusing all domains and splitting into 5 * 69 classification tasks.
A.6 Experimental settings
A.6.1 Comparison Experiments (Table 1)
@@
Our selection of baselines in this work aims to encompass a wide range of baseline categories, covering two of three primary categories in contemporary CL researches (i.e., replay-based, regularization-based, and architecture-based), with the replay-based methods not conforming to our rehearsal-free setting. The specific choice of baselines in each category is mainly based on performance comparison conclusions in recent works such as smith2023closer and kang2022forget. Therefore, we believe that our comparison study has included the most competitive and representative baselines.
@@
In CIFAR-100 and ImageNet-R(10 tasks) datasets with the ResNet-32 backbone network, we evaluate various baseline methods applicable to its model structure including Full Finetune, EWC, L2, PredKD+FeatKD, PackNet, HyperNet, WSN. All baselines are implemented on our own using a ResNet-32 with ‘option A’, i.e., leveraging the zero-padding shortcuts for increasing dimensions. In CIFAR-100 and ImageNet-R datasets with the ViT-B/16 backbone network, we mainly evaluate PEFT baseline methods including L2P, DualPrompt, CODA-Prompt, HiDe-Prompt, InfLoRA, and SD-LoRA. We cited the results of most baselines from previous works [wang2023hierarchical, wu2025sd]. For ResNet, the experiments on CIFAR-100 are conducted on NVIDIA GeForce RTX 3090 GPUs with 100 epochs of training (unless early-stop), and the ImageNet‑R experiments are carried out on NVIDIA A800 or A100 GPUs with 200 epochs of training (unless early-stop). For ViT, all experiments are conducted on NVIDIA A800 or A100 GPUs until convergence.
@@
For the EWC baseline [kirkpatrick2017overcoming], we add a regularization term to the loss function to penalize significant changes to parameters important for previously learned tasks. The importance of each parameter is estimated using the Fisher Information Matrix. The regularization coefficient $\lambda$ is set to 10, following standard practice.
@@
For the PredKD + FeatKD method [smith2023closer], we incorporate both prediction distillation and feature distillation to transfer knowledge from previous tasks to new ones. The distillation loss combines the Kullback-Leibler divergence between the soft outputs of the teacher (model trained on previous tasks) and the student (current model), as well as the mean squared error between their intermediate feature representations. The loss weights are set to $\alpha=1.0$ and $\beta=0.5$ based on preliminary tuning.
In the PackNet method [mallya2018packnet], we employ iterative pruning to allocate dedicated network weights for each task. After training on each task, we prune a certain percentage of the weights with the smallest magnitudes. Following the recommendations in the original paper, we experiment with pruning rates of 0.5, 0.75, and 0.8. We select the pruning rate of 0.8, which yields the best performance in our setting. After pruning, we fine-tune the remaining weights for an additional 10 epochs with a reduced learning rate of $1\times 10^{-4}$.
For HyperNet method [von2020continual], we primarily follow the original work in training settings, with the learning rate being 0.001, the CL loss beta being 0.05, and scheduling and transforming strategies being the same as those used by Oswald. The embedding dimension is set to 32.
For the WSN method [kang2022forget], we follow its original paper and use the default values of parameters in its official code repository. We choose the sparsity parameter $c=0.5$ which performs best as listed in the WSN literature. Other parameters are set to the following values: optimization via Adam, a learning rate initialized at 1e-3 with a minimum of 1e-6, and a patience of 6 epochs for reducing the learning rate by a factor of 2. The models are trained for 100 epochs, with a batch size of 64 for both training and testing.
For H-embedding guided hypernet, the learning rate is set to 0.0005, with the embedding loss beta and CL loss beta both set to 0.05. The scheduling and transforming strategies are set the same as those of von2020continual and the embedding dimension is also set to 32. For the learning of H-embedding, we update Eqn. 5 using gradient descent for 2000 iterations and the $f$, $g$ in HGR maximal correlation for 100 epochs respectively, both using a subset of 1000 samples from the training set.
@@
In all methods, we adhere to the principles of continual learning by not tuning hyperparameters on the full task set. In hypernet generation, special care was taken in handling batch normalization layers, especially in methods involving parameter freezing or pruning. We store and update batch normalization statistics separately for each task to ensure proper normalization during both training and inference. Also, to ensure that the parameter size of the hypernet framework should not exceed that of a normal network, we follow von2020continual and chunk the weights into size of 7000, using additional chunking embedding to generate each chunked weight.
A.6.2 Additional Comparison Experiments (Table 2)
@@
A.7 Ablation Studies
A.7.1 H-embed Hnet - LoRA (ViT) Ablation
@@
A.7.2 H-embed Hnet - Full Generation Ablation
@@
Considering the smaller data dimension and model size in this setting, the embedding size is reduced to 24 and the training iteration number is set to 5000. The backbone model on PermutedMNIST is selected to be an MLP with fully-connected layers of size 1000, 1000 as used by van2019three. We configure the learning rate as 0.0001 and the embedding loss beta as 0.05. The results are derived on NVIDIA GeForce RTX 3090 GPUs.
@@
The tasks in this setting are derived the same as in comparison studies. Yet, the backbone model is differently set to a 4-layer CNN as used by zenke2017continual. We also follow Oswald in most of the hyperparameters, configuring the learning rate to 0.0001, embedding size to 32, as well as using the same scheduling strategies. We train each method with 100 epochs and the embedding loss beta is set to 0.2 for H-embed and rand-embed hypernets. The results are derived on NVIDIA GeForce RTX 3090 GPUs.
@@
A.8 In-depth Performance Analysis